The area of a circle is 17 square meters. What is its circumference?
Answer by nerdybill(7384)

The area of a circle is 17 square meters. What is its circumference?
.
Area = (pi)R^2
where
R is radius
pi is a constant of 3.14
.
Substitute in what we know:
17 = (3.14)R^2
R^2 = 17/3.14
R^2 = 5.414
R = 2.327 meters
.
Circumference for any circle is:
circumference = (pi)D = 2R(pi)
where
D is diameter
R is radius
.
Since we now know R, substitute it in:
circumference = 2R(pi)
circumference = 2(2.327)(3.14)
circumference = 14.612 meters